# From Reasoning to Learning: A Survey on Hypothesis Discovery and Rule Learning with Large Language Models

### v1

@@ 7.2 Proactive Hypothesis Discovery
Different from relying solely on an LLM’s parametric memory to generate new evidence, Seals & Shalin (2024) propose a minimal setting for proactive hypothesis discovery. Inspired by the Wason Task from cognitive science, this task challenges LLMs to prove a formal language hypothesis of the form “if $p$ then $q$.” Here, both $p$ and $q$ are objects described in natural language, for example, “if a person is a man, then he drinks alcohol.” The task provides four cards, each with two sides representing different attributes. Initially, one side of each card is shown, displaying $p,q,\neg p,$ and $\neg q$, while the other side reveals the state of another attribute. To rigorously validate the hypothesis “if $p$ then $q$,” one must flip the $p$ card to confirm that its hidden side is $q$ (modus ponens) and flip the $\neg q$ card to check that its hidden side is $\neg p$ (modus tollens). Flipping only these two cards provides sufficient evidence for the hypothesis, while the other two cards do not offer the necessary information. Thus, in this benchmark, by proactively flipping two cards, we can determine whether the LLM can correctly identify natural language expressions of $p$ and $q$ and validate the hypothesis using a minimal action space.
@@ 7.3 Complete Loop: Real-World Hypothesis Discovery Simulation
Other works equip LLM agents with interactive environments that more closely mirror the complexity of real-world hypothesis discovery by combining planning, acting, and evidence collection. For example, Xu et al. (2023) construct a Minecraft–like world in which a “vandal” agent performs up to 26 types of actions (e.g., moving, eating, crafting) to achieve a hidden goal (such as collecting lava or crafting a particular item) and leaves behind tracks as evidence. A detective agent—driven by reinforcement learning to maximize information gain—then gathers those tracks and presents them to an LLM, which must answer a multiple-choice question about the vandal’s original objective. Because evidence collection relies on an RL policy rather than LLM planning, however, this setup evaluates only the model’s capacity to interpret evidence, not its ability to proactively generate and test hypotheses in a dynamic setting.
Building on this approach, Wang et al. (2022a) introduce 30 scientific tasks drawn from five topics in fifth-grade curricula, ranging from measuring the friction coefficient of an inclined plane to testing electrical conductivity. Here, agents must execute long action sequences and apply deductive reasoning grounded in established theories and definitions to complete each task. Likewise, Jansen et al. (2024) design 120 experiments across eight subjects (e.g., Chemistry, Archaeology), each with three difficulty levels, and allow 14 coarse-grained actions (such as “take,” “put,” and “move”). Agents are evaluated on (1) task completion, (2) execution of key experimental steps, and (3) accurate hypothesis discovery compared to a ground truth. While these virtual labs simulate multi-step procedures and test hypothesis application, their restricted action spaces support only qualitative inference and preclude the fine-grained interventions needed for quantitative rule-learning.
To address these limitations, He et al. (2024) propose puzzle environments in which agents can input arbitrary integers or letters and receive tailored feedback based on a hidden rule. In this framework, an LLM must iteratively probe the environment, uncover the underlying quantitative rule, and solve the puzzle. Performance is assessed not only by whether the agent solves the puzzle but also by human judgments of the clarity and rigor of its reasoning steps, thereby offering a finer-grained evaluation of both quantitative hypothesis generation and the quality of the model’s deductive process.

### v2

@@ 7.2 Proactive Hypothesis Discovery
Different from relying solely on an LLM’s parametric memory to generate new evidence, Seals & Shalin (2024) propose a minimal setting for proactive hypothesis discovery. Inspired by the Wason Task from cognitive science, this task challenges LLMs to prove a formal language hypothesis of the form “if $p$ then qq.” Here, both $p$ and $q$ are objects described in natural language, for example, “if a person is a man, then he drinks alcohol.” The task provides four cards, each with two sides representing different attributes. Initially, one side of each card is shown, displaying $p,q,\neg p,$ and $\neg q$, while the other side reveals the state of another attribute. To rigorously validate the hypothesis “if $p$ then qq,” one must flip the $p$ card to confirm that its hidden side is $q$ (modus ponens) and flip the $\neg q$ card to check that its hidden side is $\neg p$ (modus tollens). Flipping only these two cards provides sufficient evidence for the hypothesis, while the other two cards do not offer the necessary information. Thus, in this benchmark, by proactively flipping two cards, we can determine whether the LLM can correctly identify natural language expressions of $p$ and $q$ and validate the hypothesis using a minimal action space.
@@ 7.3 Complete Loop: Real-World Hypothesis Discovery Simulation
Other works equip LLM agents with interactive environments that more closely mirror the complexity of real‐world hypothesis discovery by combining planning, acting, and evidence collection. For example, Xu et al. (2023) construct a Minecraft–like world in which a “vandal” agent performs up to 26 types of actions (e.g., moving, eating, crafting) to achieve a hidden goal (such as collecting lava or crafting a particular item) and leaves behind tracks as evidence. A detective agent—driven by reinforcement learning to maximize information gain—then gathers those tracks and presents them to an LLM, which must answer a multiple‐choice question about the vandal’s original objective. Because evidence collection relies on an RL policy rather than LLM planning, however, this setup evaluates only the model’s capacity to interpret evidence, not its ability to proactively generate and test hypotheses in a dynamic setting.
Building on this approach, Wang et al. (2022a) introduce 30 scientific tasks drawn from five topics in fifth‐grade curricula, ranging from measuring the friction coefficient of an inclined plane to testing electrical conductivity. Here, agents must execute long action sequences and apply deductive reasoning grounded in established theories and definitions to complete each task. Likewise, Jansen et al. (2024) design 120 experiments across eight subjects (e.g., Chemistry, Archaeology), each with three difficulty levels, and allow 14 coarse‐grained actions (such as “take,” “put,” and “move”). Agents are evaluated on (1) task completion, (2) execution of key experimental steps, and (3) accurate hypothesis discovery compared to a ground truth. While these virtual labs simulate multi‐step procedures and test hypothesis application, their restricted action spaces support only qualitative inference and preclude the fine‐grained interventions needed for quantitative rule‐learning.
To address these limitations, He et al. (2024) propose puzzle environments in which agents can input arbitrary integers or letters and receive tailored feedback based on a hidden rule. In this framework, an LLM must iteratively probe the environment, uncover the underlying quantitative rule, and solve the puzzle. Performance is assessed not only by whether the agent solves the puzzle but also by human judgments of the clarity and rigor of its reasoning steps, thereby offering a finer‐grained evaluation of both quantitative hypothesis generation and the quality of the model’s deductive process.